### v1

@@ A.1.4 Gaussian Data Poisoning
Motivation. The Gaussian Unlearning Score (GUS) uses the following simple fact about Gaussian random variables to devise an unlearning test: Let $\xi\sim\mathcal{N}(0,\varepsilon_{p}^{2}\mathbb{I})$ and let $g$ be a constant with respect to $\xi$, then $\frac{\langle g,\xi\rangle}{\varepsilon_{p}\|g\|}\sim\mathcal{N}(0,1)$. In other words, if the gradient $g$ and the poison $\xi$ are statistically independent, then their normalized dot product will follow a standard normal distribution. On the other hand, when unlearning did not succeed and $g$ may depend on $\xi$, then $\frac{\langle g,\xi\rangle}{\varepsilon_{p}\|g\|}$ will deviate from a standard normal distribution. In particular, we can use the deviation of $\operatorname{\mathbb{E}}\left[\frac{\langle g,\xi\rangle}{\varepsilon_{p}\|g\|}\right]$ from 00 to measure the ineffectiveness of approximate unlearning.
For the sake of intuition, in the following, we provide an artificial example to demonstrate the change in distribution from $\mathcal{N}(0,1)$ when $\xi$ depends on $g$. Suppose the poison sample $z\in S_{\textsf{poison}}$ is generated by adding the noise $\xi_{z}$ to the base sample $(x_{\textsf{base}},y)$ in the clean training dataset. Furthermore, suppose that the gradient $g_{z}$ in the sample space w.r.t. the clean training sample $(x_{\textsf{base}},y)$ corresponding to the poison sample $z$ satisfies the relation $g_{z}=\xi_{z}$. Then, $\langle g_{z},\xi_{z}\rangle=\langle\xi_{z},\xi_{z}\rangle$ denotes a sum of $d$ many $\chi^{2}$-random variables with expectation $\varepsilon_{p}$ each, and that $\operatorname{\mathbb{E}}\left[I_{z}\right]\vcentcolon={}\operatorname{\mathbb{E}}\left[\frac{\langle g_{z},\xi_{z}\rangle}{\varepsilon_{p}\|g_{z}\|}\right]\approx\sqrt{\frac{d}{2}}$. On the other hand, when $g_{z}$ is independent of $\xi_{z}$ (for example for a model which has completely unlearnt the poison samples), we have that $I_{z}=\frac{\langle g_{z},\xi_{z}\rangle}{\varepsilon_{p}\|g_{z}\|}\sim\mathcal{N}(0,1)$ for each poison sample $z\in S_{\textsf{poison}}$. We can thus compare which of the two distributions does $I_{z}$ belong to by evaluating the mean $\frac{1}{\lvert S_{\textsf{poison}}\rvert}\sum_{z}\frac{\langle g_{z},\xi_{z}\rangle}{\varepsilon_{p}\|g_{z}\|}$. Informally speaking, further away is this mean from 00, more is the influence of the data poisons on the underlying models.
@@
6:     Let $I_{z}=\frac{\left\langle g_{z},\xi_{z}\right\rangle}{\varepsilon_{p}\|g_{z}\|_{2}}$ where $\xi_{z}$ denotes the noise used to generate the poison sample $z$.
@@
14:     Define $I_{z}=\frac{1}{\varepsilon_{p}}\langle g_{\textsf{initial},z},\xi_{z}\rangle$ and update $\mathcal{I}_{\textsf{poison}}=\mathcal{I}_{\textsf{poison}}\cup I_{z}$.
@@
22:      Compute (normalized) input gradient $g_{\text{{updated}},z}=\frac{\nabla_{x}\ell_{\theta_{\textsf{updated}}}(x_{\textsf{base}},y)}{\varepsilon_{p}\|\nabla_{x}\ell_{\theta_{\textsf{updated}}}(x_{\textsf{base}},y)\|)}$.
23:     Define $I^{\prime}_{z}=\frac{1}{\varepsilon_{p}}\langle g_{\textsf{updated},z},\xi_{z}\rangle$ and update ${\mathcal{I}}_{\textsf{updated}}={\mathcal{I}}_{\textsf{updated}}\cup I^{\prime}_{z}$.
@@
25:// For perfect unlearning, $\hat{\mu}_{\text{{updated}}}\sim\mathcal{N}(0,\nicefrac{{1}}{{P}})$. Thus, when $\hat{\mu}_{\text{{updated}}}$ is comparable to $\hat{\mu}_{\text{{initial}}}>0$ then unlearning did not succeed. //
@@
The hyperparameters used to compute the Gaussian poisons in our experiments are:
@@
Further details on the Gaussian poison attack. As we have clarified in the main text, the Gaussian poisoning attack attempts to induce a dependence between the gradient with respect to the updated model evaluated at the clean image, and the poisons $\{\xi_{z}\}_{z\in S_{\textsf{poison}}}$. Larger absolute values of this dependence statistic $\{I_{z}\}$ after unlearning, are evidence that the unlearning algorithm did not fully remove the impact of the poisons.
Interpreting the Gaussian poison attack as a membership inference attack. Consider a routine that samples a point $z$ from $\frac{1}{2}\mathcal{I}_{\textsf{poison}}+\frac{1}{2}\mathcal{I}_{\textsf{indep}}$, computes $|I_{z}|$ using the unlearned model, and then guesses that $z\in\mathcal{I}_{\textsf{poison}}$ if $|I_{z}|>\tau$. Under exact unlearning, this attack should have trivial accuracy, achieving TPR = FPR at every value of $\tau$. To illustrate, consider the right most panel from Figure 6 where unlearning is not exact since the blue histogram deviates from the teal $\mathcal{N}(0,1)$ distribution curve which represents perfect unlearning. Hence, we measure unlearning error, by the extent to which a classifier achieves nontrivial accuracy when deciding whether samples are from $\mathcal{I}_{\textsf{poison}}$ or $\mathcal{I}_{\textsf{indep}}$, in particular focusing on the true positive rate (TPR) at false positive rates (FPR) at or below 0.01 (denoted as TPR@@@@FPR=0.01). This measure corresponds to the orange bars we report in Figure 2.
@@
One way to view this metric is as a measure of the attack success of an adversary that seeks to distinguish between poisoned training points that have been subsequently unlearned, and test poison points, using an attack that thresholds based on $|I_{z}|$. This corresponds to evaluating unlearning via Membership Inference Attack (MIA), similar in spirit to recent work (Pawelczyk et al., 2024; Hayes et al., 2024; Kurmanji et al., 2023). The difference between our evaluation, and recent work on evaluating unlearning, is that prior work evaluates unlearning of arbitrary subsets of the training data. As a result, building an accurate attack requires sophisticated techniques that typically involve an expensive process of training additional models called shadow models, using them to estimate distributions on the loss of unlearned points, and then thresholding based on a likelihood ratio. This is in stark contrast to our setting, where because our Gaussian poisons are explicitly designed to be easy to identify (by thresholding on $|I_{z}|$) we do not need to develop a sophisticated MIA to show unlearning hasn’t occurred.
@@
To assess how good unlearning works, we consider how much information the Gaussian poisons leak from the model when no unlearning is performed, labeled as No unlearning in all figures. It represents the TPR at low FPR of the poisoned model before unlearning (solid orange lines in Figures 2 and 3). We evaluate the success of the unlearning process by determining if the forget set is effectively removed and if the model’s original behavior is restored. Ideally, the the TPR at low FPR should equal the FPR (dashed orange lines in Figure 2).
A.2 Unlearning Algorithms
A.2.1 Gradient Descent (GD)

### v2

@@ Appendix B Gaussian Data Poisoning
B.1 Motivation
Evaluating Gaussian poisons. The effect of data poisoning on a model $\theta$ is measured by the dependence of the model on the added perturbations $\{\xi_{z}\}_{z\in S_{\textsf{poison}}}$. Let $\theta$ be a model to be evaluated (which may or may not have been corrupted using poisons). In order to evaluate the effect of poison samples on $\theta$, for every poison sample $z\in S_{\textsf{poison}}$, we compute the normalized inner product $I_{z}=\nicefrac{{\left\langle g_{z},\xi_{z}\right\rangle}}{{\epsilon_{p}\|g_{z}\|_{2}}}\text{ with }g_{z}=\nabla_{x}\ell(\theta,(x_{\textsf{base}},y))$, where $g_{z}\in\mathbb{R}^{d}$ denotes the gradient of the model $\theta$ w.r.t. the input space $x$ when evaluated at the clean base image $(x_{\textsf{base}},y)$ corresponding to the poisoned sample $z$, and define the set $\mathcal{I}_{\textsf{poison}}=\{I_{z}\}_{z\in S_{\textsf{poison}}}$.
Interpreting the Gaussian poison attack as a membership inference attack. Consider a routine that samples a point $z$ from $\frac{1}{2}\mathcal{I}_{\textsf{poison}}+\frac{1}{2}\mathcal{I}_{\textsf{indep}}$, computes $I_{z}$ using the unlearned model, and then guesses that $z\in\mathcal{I}_{\textsf{poison}}$ if $I_{z}>\tau$. Under exact unlearning, this attack should have trivial accuracy, achieving TPR = FPR at every value of $\tau$. To illustrate, consider the right most panel from Figure 6 where unlearning is not exact since the blue histogram deviates from the teal $\mathcal{N}(0,1)$ distribution curve which represents perfect unlearning. Hence, we measure unlearning error, by the extent to which a classifier achieves nontrivial accuracy when deciding whether samples are from $\mathcal{I}_{\textsf{poison}}$ or $\mathcal{I}_{\textsf{indep}}$, in particular focusing on the true positive rate (TPR) at false positive rates (FPR) at or below 0.01 (denoted as TPR@@@@FPR=0.01). This measure corresponds to the orange bars we report in Figure 3.
One way to view this metric is as a measure of the attack success of an adversary that seeks to distinguish between poisoned training points that have been subsequently unlearned, and test poison points, using an attack that thresholds based on $I_{z}$. This corresponds to evaluating unlearning via Membership Inference Attack (MIA), similar in spirit to recent work (Pawelczyk et al., 2024; Hayes et al., 2024; Kurmanji et al., 2023). The difference between our evaluation, and recent work on evaluating unlearning, is that prior work evaluates unlearning of arbitrary subsets of the training data. As a result, building an accurate attack requires sophisticated techniques that typically involve an expensive process of training additional models called shadow models, using them to estimate distributions on the loss of unlearned points, and then thresholding based on a likelihood ratio. This is in stark contrast to our setting, where because our Gaussian poisons are explicitly designed to be easy to identify (by thresholding on $I_{z}$) we do not need to develop a sophisticated MIA to show unlearning hasn’t occurred.
To assess how good unlearning works, we consider how much information the Gaussian poisons leak from the model when no unlearning is performed, labeled as No unlearning in all figures. It represents the TPR at low FPR of the poisoned model before unlearning (solid orange lines in Figures 3 and 4). We evaluate the success of the unlearning process by determining if the forget set is effectively removed and if the model’s original behavior is restored. Ideally, the TPR at low FPR should equal the FPR (dashed black lines in Figure 3).
B.2 Gaussian Poisoning as a Hypothesis Testing Problem.
We can translate the above reasoning into membership hypothesis test of the following form:
  • $H_{0}$:
    The model $f$ was trained on $S_{\textsf{train}}$ without $\xi$ (perfect unlearning / $\xi$ is a test poison);
  • $H_{1}$:
    The model $f$ was trained on $S_{\textsf{train}}$ with $\xi$ (imperfect unlearning / no unlearning).
@@ B.1 Motivation
Constructing the test statistic. The Gaussian Unlearning Score (GUS) uses the following simple fact about Gaussian random variables to devise an unlearning test: Let $\xi\sim\mathcal{N}(0,\varepsilon_{p}^{2}\mathbb{I})$ and let $g$ be a constant with respect to $\xi$, then $\frac{\langle g,\xi\rangle}{\epsilon_{p}\|g\|}\sim\mathcal{N}(0,1)$.
  • $H_{0}$:
    Consider the model’s gradient at an base image. If the gradient $g$ is constant with respect to $\xi$, then their normalized dot product will follow a standard normal distribution.
  • $H_{1}$:
    When unlearning did not succeed and $g$ may depend on $\xi$, then $\frac{\langle g,\xi\rangle}{\epsilon_{p}\|g\|}$ will deviate from a standard normal distribution. In particular, we can use the deviation of $\operatorname{\mathbb{E}}\left[\frac{\langle g,\xi\rangle}{\epsilon_{p}\|g\|}\right]$ from 00 to measure the ineffectiveness of approximate unlearning.
B.3 An Illustrative Edge Case
Blessing of Dimensionality: Higher Input Dimension Contributes to Higher Power. Here our goal is to understand the factors that determine the success of the Gaussian poisoning method. For the sake of intuition, in the following, we provide an artificial example to demonstrate the change in distribution from $\mathcal{N}(0,1)$ when $\xi$ does not depend on $g$ to the distribution under the alternative hypothesis when $g$ depends on $\xi$. Suppose the poison sample $z\in S_{\textsf{poison}}$ is generated by adding the noise $\xi_{z}$ to the base sample $(x_{\textsf{base}},y)$ in the clean training dataset. For illustrative purposes, we will consider an extreme case.
  • $H_{0}$:
    When $g_{z}$ is constant wrt to $\xi_{z}$ (for example for a model which has completely unlearned the poison samples), we have that $I_{z}=\frac{\langle g_{z},\xi_{z}\rangle}{\epsilon_{p}\|g_{z}\|}\sim\mathcal{N}(0,1)$ for each poison sample $z\in S_{\textsf{poison}}$.
  • $H_{1}$:
    Suppose that the gradient $g_{z}$ in the sample space w.r.t. the clean training sample $(x_{\textsf{base}},y)$ corresponding to the poison sample $z$ only memorizes the poison and hence satisfies the relation $g_{z}=\xi_{z}$. Then, $\langle g_{z},\xi_{z}\rangle=\langle\xi_{z},\xi_{z}\rangle$ denotes a sum of $d$ many $\chi^{2}$-random variables with expectation $d$ and variance $2d$ and then $\operatorname{\mathbb{E}}\left[I_{z}\right]\vcentcolon={}\operatorname{\mathbb{E}}\left[\frac{\langle\xi_{z},\xi_{z}\rangle}{\sqrt{2d}}\right]=\sqrt{\nicefrac{{d}}{{2}}}$. Further, by the Central Limit Theorem, $I_{z}$ converges in distribution to $\mathcal{N}(\sqrt{\nicefrac{{d}}{{2}}},1)$.
For this special case, we thus see that our hypothesis testing problem boils down to comparing two normal distributions to each other; one with mean 00 and standard deviation $1$, and the other with mean $\mu\geq 0$ and standard deviation $1$. By the Neyman-Pearson Lemma, we know that the best true positive rate (TPR) at a given false positive rate (FPR) for this problem is given by:
$\text{TPR}(\text{FPR})=1-\Phi(\Phi^{-1}(1-\text{FPR})-\mu)\text{ with }\mu=\sqrt{\nicefrac{{d}}{{2}}}.$ (3)
This suggests that we will be able to better distinguish perfect unlearning from unlearning failure the higher the input dimension is.
Moving beyond the Edge Case. In more practical use cases, where closed-form computation is not feasible, $\mu$ likely depends on the model architecture, the optimizer, and the training dynamics. Our simple theory predicts that the Gaussian Unlearning Score should grow slower than linearly, with diminishing benefits from additional dimensions. In Figure 8, we verify that higher input dimensions make Gaussian data poisoning more effective in more complex scenarios. We trained ResNet18 models for 100 epochs on the CIFAR-10 dataset, varying the input size from 26x26x3 to 32x32x3, and subsequently unlearn 1000 data points using NGD with noise level $\sigma^{2}_{\text{NGD}}=1e-07$. The figure demonstrates the diminishing benefits of adding additional data dimensions.
B.4 Algorithms
Computing GUS. In practice, we can thus compare which of the two distributions does $I_{z}$ belong to by evaluating the mean $\frac{1}{\lvert S_{\textsf{poison}}\rvert}\sum_{z}\frac{\langle g_{z},\xi_{z}\rangle}{\epsilon_{p}\|g_{z}\|}$. Informally speaking, the further away this mean is from 00, the higher is the influence of the data poisons on the underlying models. Algorithm 1 shows how we compute GUS.
@@ B.4 Algorithms
6:     Let $I_{z}=\frac{\left\langle g_{z},\xi_{z}\right\rangle}{\epsilon_{p}\|g_{z}\|_{2}}$ where $\xi_{z}$ denotes the noise used to generate the poison sample $z$.
@@
14:     Define $I_{z}=\frac{1}{\epsilon_{p}}\langle g_{\textsf{initial},z},\xi_{z}\rangle$ and update $\mathcal{I}_{\textsf{poison}}=\mathcal{I}_{\textsf{poison}}\cup I_{z}$.
@@
22:      Compute (normalized) input gradient $g_{\text{{updated}},z}=\frac{\nabla_{x}\ell_{\theta_{\textsf{updated}}}(x_{\textsf{base}},y)}{\epsilon_{p}\|\nabla_{x}\ell_{\theta_{\textsf{updated}}}(x_{\textsf{base}},y)\|)}$.
23:     Define $I^{\prime}_{z}=\frac{1}{\epsilon_{p}}\langle g_{\textsf{updated},z},\xi_{z}\rangle$ and update ${\mathcal{I}}_{\textsf{updated}}={\mathcal{I}}_{\textsf{updated}}\cup I^{\prime}_{z}$.
@@
25:// For perfect unlearning, $\hat{\mu}_{\text{{updated}}}\sim\mathcal{N}(0,1)$. Thus, when $\hat{\mu}_{\text{{updated}}}$ is comparable to $\hat{\mu}_{\text{{initial}}}>0$ then unlearning did not succeed. //
@@
Further details on the Gaussian poison attack. As we have clarified in the main text, the Gaussian poisoning attack attempts to induce a dependence between the gradient with respect to the updated model evaluated at the clean image, and the poisons $\{\xi_{z}\}_{z\in S_{\textsf{poison}}}$. Larger values of this dependence statistic $\{I_{z}\}$ after unlearning, are evidence that the unlearning algorithm did not fully remove the impact of the poisons.
[…]
Why is the bias negative in the above? I would expect the other way around. The hyperparameters used to compute the Gaussian poisons in our experiments are:
@@
[…]
Appendix C Unlearning Evaluation Methods: Methodological Comparisons
Adversary Specific Training Data Knowledge Training Algorithm Model Architecture Trigger Required Targeted Data Poisoning $\checkmark$ $\checkmark$ $\checkmark$ × Indiscriminate Data Poisoning $\checkmark$ $\checkmark$ $\checkmark$ × Backdoor Data Poisoning $\checkmark$ × × $\checkmark$ Gaussian Data Poisoning (Ours) × × × ×
Table 3: Gaussian data poisoning has minimal knowledge requirements. Information that adversary needs to implement the corresponding data poisoning attacks considered in this work.
Adversary Image Data Language Data Tabular Data Targeted Data Poisoning $\checkmark$ × $\checkmark$ Indiscriminate Data Poisoning $\checkmark$ × $\checkmark$ Backdoor Data Poisoning × $\checkmark$ × Gaussian Data Poisoning (Ours) $\checkmark$ $\checkmark$ $\checkmark$
Table 4: Gaussian data poisoning is compatible with all common data formats. Data compatibility of different data poisoning methods considered in this work.
Adversary Resnet18 GPT2 Targeted Data Poisoning $\approx\nicefrac{{1}}{{2}}$ hour N/A Indiscriminate Data Poisoning $\approx 1$ week N/A Backdoor Data Poisoning N/A $\approx 7$ minutes Gaussian Data Poisoning (Ours) $\approx\nicefrac{{1}}{{2}}$ minute $\approx 1$ minute
Table 5: Gaussian data poisoning is more computationally efficient.
Our Gaussian data poisoning method overcomes key limitations of the other data poisoning-based unlearning evaluations through four critical dimensions:
  1. $\bullet$
    Computational Efficiency: Gaussian poisoning offers a dramatic improvement in computational complexity compared to any other unlearning evaluation (see Table 5). Unlike existing methods that require complex optimization procedures taking minutes (e.g., targeted and indiscriminate data poisoning attacks), our approach involves a simple gradient computation and dot product calculation, completing in mere seconds. We include a comprehensive runtime comparison demonstrating this efficiency across different model architectures and datasets.
  2. $\bullet$
    Minimal Knowledge Requirements: Gaussian poisoning stands out for its minimal prerequisite knowledge. Where other unlearning evaluation methods demand extensive model or dataset information, our approach requires minimal contextual understanding. Table 3 provides a comprehensive comparison illustrating the knowledge constraints of different unlearning evaluation techniques.
  3. $\bullet$
    Data Compatibility: Unlike existing targeted attacks limited to specific domains, Gaussian poisoning demonstrates remarkable versatility. Our method successfully operates across diverse data types including text, images, and tabular data (see Table 4. This generalizability is particularly significant given the challenges of extending existing methods to emerging domains like large language models.
  4. $\bullet$
    Privacy Impact Measurement: Crucially, Gaussian poisoning directly addresses privacy concerns by precisely measuring individual sample information retention. Existing unlearning evaluations and data poisoning methods fail to provide this granular privacy impact assessment, making our approach uniquely valuable for understanding true unlearning effectiveness.
Appendix D Implementation Details
D.1 Existing Data Poisoning Attacks
We provide the key implementation details below:
  1. $\bullet$
[…]
  2. $\bullet$
D.1.1 Targeted Data Poisoning for Image Classification
$\min_{\Delta\in\Gamma}\ell\left(f(x_{\textsf{target}},\theta(\Delta)),y_{\mathrm{adv}}\right)\quad\text{where}$ (4)
where the constraint set $\Gamma\vcentcolon={}\left\{\Delta\mid{}\|\Delta(x)\|_{\infty}\leq\epsilon_{p}\forall x\in S_{\textsf{poison}}\right\}.$ However, since directly solving (5) is computationally intractable due to its bi-level nature, Geiping et al. (2021) has opted for the approach to implicitly minimize the adversarial loss such that for any model $\theta$,
$\nabla_{\theta}(\ell(f(x_{\textsf{target}},\theta),y_{\textsf{advs}}))\approx\frac{\sum_{i=1}^{P}\nabla_{\theta}\ell\left(f(x^{i}+\Delta^{i},\theta),y^{i}\right)}{P}..$ (6)
1: Clean network $f(\cdot;\theta_{\mathrm{clean}})$ trained on uncorrupted base images $S_{\textsf{clean}}$ The target $(x_{\textsf{target}},y_{\textsf{target}})$ and the adversarial label $y_{\textsf{advs}}$ Poison budget $P$ and perturbation bound $\epsilon_{p}$ Number of restarts $R$ and optimization steps $M$
3:for $r=1,\dots R$   restarts do
4:     Randomly initialize perturbations $\Delta$ s.t. $\|\Delta\|_{\infty}\leq\epsilon_{p}$.
5:     for $k=1,\dots,M$   optimization steps do
  • Clean dataset $S_{\textsf{clean}}$ is the CIFAR-10 training set;
  • Set a Perturbation bound $\epsilon_{p}$ of 16.
  • Generate $\Delta$ using the algorithm outlined in Algorithm 3
D.1.2 Backdoor Data Poisoning for Language Sentiment Analysis
  • Clean dataset $S_{\textsf{clean}}$ is the IMDb reviews training set;
  • Set the maximum sequence length of the tokenizer to 128.
D.1.3 Indiscriminate Data Poisoning
For a given poison budget $b_{p}$ and perturbation bound $\epsilon_{p}$, we generate the poison samples by following the Gradient Canceling (GC) procedure of Lu et al. (2023; 2024), a state-of-the-art indiscriminate poisoning attack in machine learning. In Gradient Canceling (GC) procedure, the adversary first finds a bad model $\theta_{\mathrm{low}}$ that has low-performance accuracy on the test dataset and then computes the perturbations $\Delta$ by solving the minimization problem
$\operatorname*{argmin}_{\Delta\in\Gamma}~{}\tfrac{1}{2}\|\widehat{\operatorname{\mathbb{E}}}_{(x,y)\in S_{\textsf{clean}}}\left[\nabla_{\theta}\ell((x,y);\theta_{\mathrm{low}})\right]+\widehat{\operatorname{\mathbb{E}}}_{(x,y)\in S_{\textsf{poison}}}\left[\nabla_{\theta}\ell((x+\Delta(x),y_{i});\theta_{\mathrm{low}}\right])\|_{2}^{2},$ (8)
where the constraint set $\Gamma\vcentcolon={}\left\{\Delta\mid{}\|\Delta(x)\|_{\infty}\leq\epsilon_{p}\forall x\in S_{\textsf{poison}}\right\}.$ Informally speaking, the above objective function enforces that the generated poison points are such that $\theta_{\mathrm{low}}$ has vanishing (sub)gradients over the corrupted training dataset, and is thus close to a local minimizer of the training objective using the corrupted dataset. The model $\theta_{\mathrm{low}}$ is generated by the GradPC procedure of Sun et al. (2020), which is a gradient-based approach to finding a set of corrupted parameters that returns the lowest test accuracy within a certain distance from an input trained parameter. We provide the pseudocode of this attack in Algorithm 4.
Algorithm 4 Gradient Canceling (GC) Attack (Lu et al., 2023)
1: An uncorrupted clean dataset $S_{\textsf{clean}}$ Target network $f(\cdot;\theta_{\mathrm{low}})$ generated by GradPC (Sun et al., 2020) Poisoning budget $b_{p}$ and perturbation bound $\epsilon_{p}$ Step size $\eta$
4:for $t=1,2,\dots$ do
6:     Calculate loss $\mathcal{L}=~{}\tfrac{1}{2}\|{g}_{c}+{g}_{p}\|_{2}^{2}$.
  • Clean dataset $S_{\textsf{clean}}$ is the CIFAR-10 training set;
  • Perturbation bound $\epsilon_{p}$ is set to be infinite. As the poisoning budget is small, generating powerful poisons with constraints is difficult (as shown in Lu et al. (2023)). Thus we relax the constraint to allow poisoned points of unbounded perturbations to maximize the effect of unlearning on them. Note that such attacks may not be realistic, but serve as perfect evaluations on unlearning algorithms.
  • Target parameters $\theta_{\mathrm{low}}$ are generated by GradPC with a budget of $\epsilon_{w}=1$, where $\epsilon_{w}$ measures the L2 distance between $\theta_{\mathrm{low}}$ and the clean parameter.
D.2 Unlearning Algorithms
D.2.1 Gradient Descent (GD)